# Determinant of the OU matrix of a braid diagram

### v1

@@ 1 Introduction
A braid is a set of strands attached to two parallel horizontal bars in $\mathbb{R}^{3}$, where each strand runs from one bar to the other bar without returning. We say a braid is an $n$-braid when the braid is consisting of $n$ strands. A braid diagram is a diagram of a braid depicted on $\mathbb{R}^{2}$. Each braid diagram is represented by a word consisting of $\sigma_{i}$ and ${\sigma_{i}}^{-1}$, where $\sigma_{i}$ (resp. ${\sigma_{i}}^{-1}$) stands for a positive (resp. negative) crossing, as shown in Figure 1.
@@
The warping degree has been defined for link diagrams and studied to measure a complexity of link diagrams in a sense of over/under information (see, for example, [5, 6, 10]). Considering the order of strands, the warping degree of a braid diagram [^1] In [8], a labeling named “warping labeling” was also defined for twisted virtual braid diagrams from a different point of view of the warping degree, following a rule of the warping degree of the closure. was also defined in [11] as follows. Let $s_{1},s_{2},\dots,s_{n}$ be the strands of an $n$-braid diagram $B$ which are positioned in the order from the left-hand side to the right-hand side at the top of the braid diagram, as shown in Figure 2. The warping degree of $(s_{i},s_{j})$, denoted by $\mathrm{wd}(s_{i},s_{j})$, is the number of crossings such that $s_{i}$ is under $s_{j}$. Let $\mathbf{s}$ be a sequence of the strands in some order. We denote the pair of $B$ and $\mathbf{s}$ by $B_{\mathbf{s}}$. The warping degree of $B_{\mathbf{s}}$, denoted by $\mathrm{wd}(B_{\mathbf{s}})$, is the total number of crossings such that $s_{i}$ is under $s_{j}$ where $s_{i}$ is positioned ahead of $s_{j}$ in the sequence $\mathbf{s}$. The warping degree of a braid diagram $B$, denoted by $\mathrm{wd}(B)$, is the minimum value of $\mathrm{wd}(B_{\mathbf{s}})$ for all $\mathbf{s}$.
@@
The braid diagram $B$ in Figure 2 has $\mathrm{wd}(s_{1},s_{2})=1$, $\mathrm{wd}(s_{2},s_{1})=0$. For the sequences of strands $\mathbf{s}=(s_{1},s_{2},s_{3},s_{4},s_{5})$ and $\mathbf{s}^{\prime}=(s_{2},s_{3},s_{1},s_{4},s_{5})$, $B$ has $\mathrm{wd}(B_{\mathbf{s}})=3$ and $\mathrm{wd}(B_{\mathbf{s}^{\prime}})=1$, respectively. We can determine the warping degree to be $\mathrm{wd}(B)=1$ by checking the the warping degrees for all the $5!$ orders or applying Theorem 2.
@@
When the set of strands $S=\{s_{1},s_{2},\dots,s_{n}\}$ of an $n$-braid diagram $B$ can be divided as $S=S_{1}\oplus S_{2}\oplus\dots\oplus S_{k}$ with $k$ non-empty sets $S_{i}$ $(k\geq 2)$ so that $\mathrm{wd}(s_{i},s_{j})=0$ for any $s_{i}\in S_{l}$ and $s_{j}\in S_{m}$ for $l<m$, we say $B$ is layered. We call each braid diagram $B_{i}$ consisting of the strands in $S_{i}$ a layer and denote $B=B_{1}\oplus B_{2}\oplus\dots\oplus B_{k}$ (see Figure 3). In particular, we say an $n$-braid diagram $B$ is completely layered when $B$ is layered with $n$ layers where each layer is consisting of a single strand. By definition, $B$ is completely layered if and only if $\mathrm{wd}(B)=0$. Therefore, the warping degree represents how far a braid diagram $B$ is from a completely layered braid diagram, like the warping degree of a link diagram $D$ represents how far $D$ is from a monotone diagram, which is a diagram of a trivial link.
Figure 3: A layered diagram $B$ with the layers $B_{1}$ and $B_{2}$ of the sets of strands $S_{1}=\{s_{1},s_{3},s_{4}\}$ and $S_{2}=\{s_{2},s_{5}\}$. The determinants are $\mathrm{det}(B)=-6$, $\mathrm{det}(B_{1})=3$, and $\mathrm{det}(B_{2})=-2$.
The warping degree of a braid diagram is useful to measure a complexity of the closure as well. For a closure $D$ of a pure braid $B$, we have the inequality [^2] The non-pure version of the inequality was also shown in [11] with some condition of the sequence of strands. $u(D)\leq\mathrm{wd}(B)$ shown in [11], namely the warping degree is an upper bound of the unlinking number $u(D)$ of the link diagram $D$. Even for non-pure braids, the warping degree has been used to estimate the unknotting number of the closure for weaving knots in [11]. However, finding the warping degree is not always easy for large number of strands; each $n$-braid has $n!$ orders of strands. We call the problem to find the value of the warping degree the warping degree problem.
@@
Let $B$ be a layered braid diagram with layers $B_{1}$, $B_{2}$. Then, the determinant of the OU matrix of $B$ is the product of the determinants of $B_{1}$ and $B_{2}$, namely,
@@
The rest of the paper is organized as follows. In Section 2, we define and study the OU matrix and prove Theorem 1. In Section 3, we discuss the warping degree problem in terms of the linear ordering problem on the OU matrices, and prove Theorem 2. In Section 4, diagram-independent invariants of positive braids are introduced as an application with examples.
@@ 2.3 Properties of OU matrix
Proof of Lemma 2.  On $M_{B}(\pi)$, each $(i,j)$-component represents the number of crossings where $s_{\pi(i)}$ is under $s_{\pi(j)}$, which appears at the $(n+1-i,n+1-j)$-component in $M_{B}({\pi}^{\prime})$ since the $i$th component of $\pi$ coincides with the $(n+1-i)$th component of ${\pi}^{\prime}$. ∎
@@
The $\pi(i)$th strand of $BC$ is the $\pi(i)$th strand in $B$, and the $\pi\rho_{B}(i)$th strand in $C$. Then the $(i,j)$-component of $M_{BC}(\pi)$, which is the number of crossings where $s_{\pi(i)}$ is under $s_{\pi(j)}$, is the sum of the number of crossings where the $\pi(i)$th strand is under the $\pi(j)$th strand in $B$ and the crossings where the $\pi\rho_{B}(i)$th strand is under the $\pi\rho_{B}(j)$th strand in $C$, which is the sum of the $(i,j)$-components of $M_{B}(\pi)$ and $M_{C}(\pi\rho_{B})$. ∎
@@ 2.4 Proof of Theorem 1
Let $B$ be a layered braid diagram with two layers $B_{1}$ and $B_{2}$ of the sets of strands $S_{1}$ and $S_{2}$, respectively, where $B_{1}$ is over $B_{2}$. Take a sequence of strands $\mathbf{s}$ of $B$ so that any strand in $S_{1}$ is positioned ahead of any strand in $S_{2}$. Then, the OU matrix $M(B_{\mathbf{s}})$ consists of the four blocks as
@@
Proof of Theorem 1.  Let $B$ be a layered braid diagram with two layers $B_{1}$ and $B_{2}$ of the set of strands $S_{1}$ and $S_{2}$, respectively, where $B_{1}$ is over $B_{2}$. Take a sequence of strands $\mathbf{s}$ same to Lemma 3. By Lemma 3 and the well-known formula
@@ 3.1 Linear ordering problem
Proof of Theorem 2.  When $\mathrm{wd}(B)=0$, we have $f_{B}(\pi)=0$ for some permutation $\pi$. Then $\mathrm{det}(M_{B}(\pi))=\mathrm{det}(B)=0$ since $M_{B}(\pi)$ is a strictly upper triangular matrix. Hence, $\mathrm{wd}(B)\neq 0$ when $\mathrm{det}(B)\neq 0$
@@ 3.2 Weaving braids
In this section, we denote the canonical diagram of the braid $B_{W}(p,q)$ as the same notation $B_{W}(p,q)$. The following proposition was proved in [11] to estimate the unknotting number of weaving knots.
@@ 4.4 Positive pure braid
For the warping degree, we have the following corollary from Proposition 8.

### v2

@@ 1 Introduction
A braid is a disjoint union of strands attached to two parallel horizontal bars in $\mathbb{R}^{3}$, where each strand runs from one bar to the other bar without returning. We say a braid is an $n$-braid when the braid is consisting of $n$ strands. A braid diagram is a diagram of a braid depicted on $\mathbb{R}^{2}$. Each braid diagram is represented by a word consisting of $\sigma_{i}$ and ${\sigma_{i}}^{-1}$, where $\sigma_{i}$ (resp. ${\sigma_{i}}^{-1}$) stands for a positive (resp. negative) crossing, as shown in Figure 1.
@@
The warping degree has been defined for link diagrams and studied to measure a complexity of link diagrams in a sense of over/under information (see, for example, [5, 6, 10]). Considering the order of strands, the warping degree of a braid diagram [^1] In [8], a labeling named “warping labeling” was also defined for twisted virtual braid diagrams from a different point of view of the warping degree, following a rule of the warping degree of the closure. was also defined recently in [11] as follows. Let $s_{1},s_{2},\dots,s_{n}$ be the strands of an $n$-braid diagram $B$ which are positioned in the order from the left-hand side to the right-hand side at the top of the braid diagram, as shown in Figure 2. The warping degree of $(s_{i},s_{j})$, denoted by $\mathrm{wd}(s_{i},s_{j})$, is the number of crossings such that $s_{i}$ is under $s_{j}$. Let $\mathbf{s}$ be a sequence of the strands in some order, and we denote the pair of $B$ and $\mathbf{s}$ by $B_{\mathbf{s}}$. The warping degree of $B_{\mathbf{s}}$, denoted by $\mathrm{wd}(B_{\mathbf{s}})$, is the total number of crossings such that $s_{i}$ is under $s_{j}$ where $s_{i}$ is positioned ahead of $s_{j}$ in the sequence $\mathbf{s}$. The warping degree of a braid diagram $B$, denoted by $\mathrm{wd}(B)$, is the minimum value of $\mathrm{wd}(B_{\mathbf{s}})$ for all $\mathbf{s}$.
@@
The braid diagram $B$ in Figure 2 has $\mathrm{wd}(s_{1},s_{2})=1$, $\mathrm{wd}(s_{2},s_{1})=0$. For the sequences of strands $\mathbf{s}=(s_{1},s_{2},s_{3},s_{4},s_{5})$ and $\mathbf{s}^{\prime}=(s_{2},s_{3},s_{1},s_{4},s_{5})$, $B$ has $\mathrm{wd}(B_{\mathbf{s}})=3$ and $\mathrm{wd}(B_{\mathbf{s}^{\prime}})=1$. We can determine the warping degree to be $\mathrm{wd}(B)=1$ by checking the warping degrees for all the $5!$ orders or applying Theorem 2.
@@
When the set of strands $S=\{s_{1},s_{2},\dots,s_{n}\}$ of an $n$-braid diagram $B$ can be divided as the disjoint union $S_{1}\amalg S_{2}\amalg\dots\amalg S_{k}$ of $k$ non-empty sets $S_{1},S_{2},\dots,S_{k}$ for some integer $k\geq 2$ so that $\mathrm{wd}(s_{i},s_{j})=0$ for any $s_{i}\in S_{l}$ and $s_{j}\in S_{m}$ for $l<m$, we say $B$ is layered. For each $i\in\{1,2,\dots,k\}$, we assume the union of strands in $B$ which belong to $S_{i}$ as a #$\#(S_{i})$-braid diagram, regardless of the position of the endpoints. For example, we assume the diagrams shown in the center and the right-hand side in Figure 3 as a $3$-braid and $2$-braid diagrams, respectively. We call each braid diagram $B_{i}$ consisting of the strands in $S_{i}$ a layer and denote $B=B_{1}\oplus B_{2}\oplus\dots\oplus B_{k}$. In particular, we say an $n$-braid diagram $B$ is completely layered when $B$ is layered with $n$ layers where each layer is consisting of a single strand. By definition, $B$ is completely layered if and only if $\mathrm{wd}(B)=0$. Therefore, the warping degree represents how far a braid diagram $B$ is from a completely layered braid diagram, like the warping degree of a link diagram $D$ represents how far $D$ is from a monotone diagram, which is a diagram of a trivial link.
Figure 3: A layered diagram $B=B_{1}\oplus B_{2}$ with the layers $B_{1}$ and $B_{2}$ of the sets of strands $S_{1}=\{s_{1},s_{3},s_{4}\}$ and $S_{2}=\{s_{2},s_{5}\}$. The determinants are $\mathrm{det}(B)=-6$, $\mathrm{det}(B_{1})=3$ and $\mathrm{det}(B_{2})=-2$, and we have $\mathrm{det}(B)=\mathrm{det}(B_{1})\mathrm{det}(B_{2})$.
The warping degree of a braid diagram is useful to measure a complexity of the closure as well. For a closure $D$ of a pure braid $B$, we have the inequality [^2] The non-pure version of the inequality was also shown in [11] with some condition of the sequence of strands. $u(D)\leq\mathrm{wd}(B)$ shown in [11], namely the warping degree is an upper bound of the unlinking number $u(D)$ of the link diagram $D$. Even for non-pure braids, the warping degree was employed to estimate the unknotting number of the closure for weaving knots in [11]. However, finding the warping degree is not always easy for large number of strands; each $n$-braid has $n!$ orders of strands. We call the problem to find the value of the warping degree the warping degree problem.
@@
Let $B$ be a layered braid diagram with layers $B_{1}$, $B_{2}$. The determinant of the OU matrix of $B$ is the product of the determinants of $B_{1}$ and $B_{2}$, namely,
@@
The rest of the paper is organized as follows. In Section 2, we define and study the OU matrix and prove Theorem 1. In Section 3, we discuss the warping degree problem in terms of the “linear ordering problem” on the OU matrices, and prove Theorem 2. In Section 4, diagram-independent invariants of positive braids are introduced as an application with examples.
@@ 2.3 Properties of OU matrix
Proof of Lemma 2.  On $M_{B}(\pi)$, each $(i,j)$-component represents the number of crossings where $s_{\pi(i)}$ is over $s_{\pi(j)}$, which appears at the $(n+1-i,n+1-j)$-component in $M_{B}({\pi}^{\prime})$ since the $i$th component of $\pi$ coincides with the $(n+1-i)$th component of ${\pi}^{\prime}$. ∎
@@
The $\pi(i)$th strand of $BC$ is the $\pi(i)$th strand in $B$, and the $\pi\rho_{B}(i)$th strand in $C$. Then the $(i,j)$-component of $M_{BC}(\pi)$, which is the number of crossings where $s_{\pi(i)}$ is over $s_{\pi(j)}$, is the sum of the number of crossings where the $\pi(i)$th strand is over the $\pi(j)$th strand in $B$ and the crossings where the $\pi\rho_{B}(i)$th strand is over the $\pi\rho_{B}(j)$th strand in $C$, which is the sum of the $(i,j)$-components of $M_{B}(\pi)$ and $M_{C}(\pi\rho_{B})$. ∎
@@ 2.4 Proof of Theorem 1
Let $B=B_{1}\oplus B_{2}$ be a layered braid diagram with two layers $B_{1}$ and $B_{2}$ of the sets of strands $S_{1}$ and $S_{2}$, respectively. Take a sequence of strands $\mathbf{s}$ of $B$ so that any strand in $S_{1}$ is positioned ahead of any strand in $S_{2}$. Then, the OU matrix $M(B_{\mathbf{s}})$ consists of the four blocks as
@@
Proof of Theorem 1.  Let $B=B_{1}\oplus B_{2}$ be a layered braid diagram with two layers $B_{1}$ and $B_{2}$ of the set of strands $S_{1}$ and $S_{2}$, respectively. Take a sequence of strands $\mathbf{s}$ same to Lemma 3. By Lemma 3 and the well-known formula
@@ 3.1 Linear ordering problem
Proof of Theorem 2.  When $\mathrm{wd}(B)=0$, we have $f_{B}(\pi)=0$ for some permutation $\pi$. Then $\mathrm{det}(B)=\mathrm{det}(M_{B}(\pi))=0$ since $M_{B}(\pi)$ is a strictly upper triangular matrix. Hence, $\mathrm{wd}(B)\neq 0$ when $\mathrm{det}(B)\neq 0$
@@ 3.2 Weaving braids
In this section, we denote the canonical diagram of the braid $B_{W}(p,q)$ by the same notation $B_{W}(p,q)$. The following proposition was proved in [11] to estimate the unknotting number of weaving knots.
@@ 4.4 Positive pure braid
As for the warping degree, we have the following corollary from Proposition 8.